OpenStudy (anonymous):

√64 + √9 11 √53 35 8√3

OpenStudy (anonymous):

3^2 is 9, and 8^2 is 64 What can you do with that?

OpenStudy (anonymous):

@brill

OpenStudy (anonymous):

can you help me out i dont really know how to solve these kind of questions

OpenStudy (anonymous):

Well let me give an example \[6^2=36\] and also \[\sqrt{36}=6\]

OpenStudy (anonymous):

Does that help you understand square roots?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok so using the above logic If I tell you \[3^2=9\] What is the \[\sqrt{9}\]?

OpenStudy (anonymous):

its 3?

OpenStudy (anonymous):

Yep And so the problem now becomes \[\sqrt{64}+3\]

OpenStudy (anonymous):

And now I tell you that 8^2=64 So what is the square root of 64?

OpenStudy (anonymous):

Oh and BTW do you understand how I'm getting the squares?

OpenStudy (anonymous):

the square root of 64 is 8?

OpenStudy (anonymous):

yea i do

OpenStudy (anonymous):

@brill

OpenStudy (anonymous):

Yup so now you have 8+3= ? And that is your answer

OpenStudy (anonymous):

thank you for your help!

OpenStudy (anonymous):

No problem :)